OpenStudy (anonymous):

15x+15 ------- x^2-1

OpenStudy (anonymous):

Factor out the 15 giving you: 15(x+1) Factor x^2 -1 into: (x-1)(x+1) You should end up with: 15/(x-1)

OpenStudy (anonymous):

Are there any common factors you can take out?

OpenStudy (callisto):

\[\frac{15x+15}{x^2-1}=\frac{15(x+1)}{(x-1)(x+1)}\]Can you cancel the common factor in the denominator and numerator?